40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16237     Entry time: Fri Jul 2 12:42:56 2021
Author: Anchal, Paco, Gautam 
Type: Summary 
Category: LSC 
Subject: snap file changed for MICH 

We corrected the MICH locking snap file C1configure_MI.req and saved an updated C1configure_MI.snap. Now the 'Restore MICH' script in IFO_CONFIGURE>!MICH>Restore MICH works. The corrections included adding the correct rows of PD_DOF matrices to be at the right settings (use AS55 as error signal). The MICH_A_GAIN and MICH_B_GAIN needed to be saved as well.

We also were able to get to PRMI SB resonance. PRM was misalgined earlier from optimal position and after some manual aligning, we were able to get it to lock just by hitting IFO_CONFIGURE>!PRMI>Restore PRMI SB (3f).

ELOG V3.1.3-